# Benjamin S. Manning MIT John J. Horton

### v3

@@ 2.3 Theory as guide towards generalizability
Economic theories—and theories from other social sciences that embrace methodological individualism (Friedman, 1953)—are well-suited to be cast into prompts precisely because good theory requires that they be grounded only in the choices and actions of individuals. Specifically, we consider a prompt to be “theoretically grounded” when it instructs the LLM to make predictions about how a human would respond based on some underlying theory or model of human behavior. For example, a utility function that trades off one’s own payoff against inequality might become “You value your own earnings but dislike outcomes where you earn far more than others.” A highly capable and tool-using LLM could even be given that utility function and parameters directly (e.g., $u(x_{self},x_{other})=\ldots$). A prospect-theoretic model reasonably maps to “You are risk averse in gains, risk seeking in losses, and probability weight the very likely and very unlikely outcomes”—and of course these could be parameterized as well.
This definition is necessarily broad because what constitutes a theory is also broad. One might consider good theories to be those that are parsimonious, predictive, and interpretable. We view these as effective guides for constructing prompts that are grounded in theory. But just as there is no universal rule as to what makes a theory good, there is not a singular all-encompassing playbook for candidate prompts. Here, the researcher must leverage their expertise appropriately to determine which theories are most relevant to the domain where predictions are desired. This should be relatively straightforward in practice. One can even simply ask an LLM to construct a candidate set of agents based on a given paper and then adjust the prompts accordingly. For example, the attached notebook—which took 15 minutes to code—contains an example where we take PDFs of 5 well-known papers in the behavioral economics literature and generate 10 agents from each. [^9] https://expectedparrot.com/content/32751dae-7a69-430d-9f1a-0c3f50ccef5b Whether these agents are effective is an empirical question, but one that is easily testable.
@@ 2.4 A brief summary of the approach
    Optimize prompts on Training Data. Optimize the given prompts to best match the training data. This might involve selecting a mixture of prompts or adjusting trait parameters to minimize some statistical distance from observed responses. Confirm that the optimized sample outperforms the baseline LLM off-the-shelf on the training data.
@@ 2.5 Methods for optimizing prompts in-sample
We briefly describe two methods for optimizing prompts with a given set of training data. The first, the selection method, assumes we have a finite library of candidate prompts and selects (or mixes) them to best fit the training data. We apply this method to the experiments presented in Section 3 with data from Arad and Rubinstein (2012). An early version of this idea was suggested by Horton (2023), and several others have explored applications (Leng et al., 2024; Xie et al., 2025; Bui et al., 2025). The second, the construction method, parametrizes a prompt template with numeric trait dimensions and optimizes those parameters. To the best of our knowledge, this method is novel. An application of the method is presented in Appendix A using data from Charness and Rabin (2002).
@@
A finite set of unique candidate natural language prompts is first specified. For each prompt $\theta\in\Theta$, the LLM is used to generate a predicted distribution $\hat{P}_{\theta}$. Let $P$ represent the observed ground-truth human distribution. The objective is to solve
@@
Alternatively, a prompt template can be parameterized by numeric trait variables. This is best illustrated with an example. Suppose $\phi_{1}$ and $\phi_{2}$ capture degrees of self-interest and inequity aversion, respectively. For instance, the prompt could be:
@@
$\hat{P}_{\theta}$ denotes the distribution induced by the LLM under parameter vector $\theta=(\phi_{1},\phi_{2})$. Given an observed human distribution $P$, the optimal parameters are found by solving $\min_{\theta}\;d\bigl{(}P,\hat{P}_{\theta}\bigr{)}$. In practice, this can be solved using any derivative-free optimization algorithm, such as Bayesian optimization or evolutionary algorithms.
@@
To assess whether $\bm{\theta^{\star}}$ generalizes, we compare predictive fit against the baseline $\hat{P}_{0}$. This is especially appealing because many measures of distances between distributions (like the KL-divergence) are not easily interpretable. Furthermore, improvement over the baseline provides direct evidence that designing and optimizing prompts is a worthwhile endeavor in the first place. [^12] Strong baseline performance would simply reflect the LLM’s already high off-the-shelf predictive capacity.
Relative improvement can be computed as a difference-in-distances: $d\bigl{(}P,\hat{P}_{0}\bigr{)}\;-\;d\bigl{(}P,\hat{P}_{\bm{\theta^{\star}}}\bigr{)}$. A positive difference indicates that the prompts provide better predictive power than the baseline. When we optimize a prompt, we seek those that yield this positive difference on both the training and testing data—that would suggest that the set of prompts generalizes.
@@ 3 Predicting behavior in novel strategic games
We begin by briefly reviewing the level-$k$ model of strategic reasoning that originally motivated AR. This model provides plausible theoretical underpinnings linking the training, testing, and novel games in this section. We then describe the structure of the original 11-20 money request game and outline our procedure for constructing theory-grounded AI agents, optimizing their parameters on human data from AR’s original experiment, and validating their predictive performance on distinct but related variants also studied by AR. Finally, we introduce an entirely new set of games adapted from AR’s original design but featuring distinct numeric ranges and a novel participant sample recruited from Prolific. Agents are evaluated on their ability to predict human behavior in these never-before-seen games.
Although this section focuses on strategic-reasoning games, the same procedure can be applied in any setting. In Appendix A, we replicate the process using the allocation games from Charness and Rabin (2002), which explore social preferences. Two additional important differences in that section are that we: i) optimize across multiple related settings rather than a single training setting, and ii) utilize the construction method to build agents. Using a preregistered experiment with entirely novel human responses as the target data, we find similar improvements.
@@ 3.1 Arad and Rubinstein (2012)’s 11-20 money request game
The level-$k$ model posits that players differ in how many steps ahead they consider when forming their strategies (Nagel, 1995; Stahl and Wilson, 1994, 1995). The model defines different types of players, from level-0 to level-$k$. Level-0 players use some pre-defined arbitrary decision rule, while level-$k$ players ($k\geq 1$) best respond assuming others are level-(k1)(k-1) reasoners. Such a model highlights the idea that players’ behavior depends not only on their decisions but also on their beliefs about how other people think.
@@
Interestingly, the 11-20 game does not have a pure strategy Nash equilibrium. The best response to any choice greater than 11 is to undercut the opponent by 1. But if the other player chooses 11, also selecting 11 is strictly dominated by every other strategy. The top row of Table 1 shows the unique symmetric mixed strategy Nash equilibrium for the game, with most of its density lying between 15-17 (levels 3 to 5).
@@
Finally, Table 1 provides results from two additional variants of the game from AR. Both variants of the game still involve two players selecting numbers between 11 and 20. They differ from the basic version in their payouts (see Appendix B for the full instructions). In the cycle version, players can earn a bonus of 20 shekels by undercutting the other’s request by exactly one or by selecting 20 when the other selects 11. This version has the same unique mixed-strategy Nash Equilibrium as the basic game because the only affected strategy profiles are (20,11) and (11,20), which are outside the equilibrium support. However, if there is a distribution of level-$k$ thinkers—as opposed to perfectly rational game-theoretic best-responders—such a change might significantly impact play.
The costless version has the same bonus structure as the basic game, but with a different payout for choices below 20. Here, requesting 20 yields 20 shekels outright, while choosing a lower amount guarantees 17 shekels plus a 20-shekel bonus if the lower request is exactly one less than the other player’s. It is comparatively “costless” to continue undercutting. This game induces a symmetric mixed strategy Nash equilibrium that is more uniform across the choice set than the basic versions.
The human responses from these game variants, also from a similar sample of college students, are noticeably shifted towards 18-20 shekels—even with the basic and cycle sharing an equilibrium. AR attributes this to the increased salience and payoff of selecting a higher number. AR concluded that the collective results from these three experiments are best explained by a mix of strategic types consisting of level-0, level-1, level-2, level-3, and random choosers.
@@ 3.2 Optimizing AI agents in-sample
We begin by eliciting the baseline AI’s response distribution ($\hat{P}_{0}$). We prompt Gpt-4o to play the basic version of the game 1,000 times without any additional instructions, setting the temperature to 1. Figure 1 displays the results. The left panel shows the empirical PMF of the baseline AI responses (red), along with the empirical human distribution $P$ from AR’s original experiment (vertical black lines). The baseline AI almost exclusively selects 19 shekels (87%), demonstrating limited variability. Using the forward KL-divergence as $d(\cdot,\cdot)$ with the human distribution as the reference, the difference between these distributions is $d(P,\hat{P}_{0})=2.7$.
@@
Notes: This figure displays empirical PMFs for three samples playing the basic 11-20 money request game: human subjects from Arad and Rubinstein (the vertical black lines in both panels), the off-the-shelf baseline (left panel), and responses from our selected AI agents based on the weights (right panel).
@@
We elicit response distributions $\hat{P}_{\theta}$ for each candidate prompt $\theta\in\Theta_{Strategic}$ by prompting Gpt-4o 100 times per prompt. [^13] We then elicit each AI agent’s responses using Chain-of-Thought prompting, which encourages step-by-step reasoning before producing a final answer (Wei et al., 2024). We implement this through two sequential prompts. Prompt 1: {11-20 game instructions}. Reason out a few settings according to your personality and how others might respond. Prompt 2: {11-20 game instructions}. You previously had the following thoughts: {Response to prompt 1}. What amount of money would you request?. This procedure creates the agents (and their response distributions) over which we then optimize. We then employ the selection method to identify the optimal mixture of prompts ($\hat{P}_{\theta}$’s) that minimizes the absolute difference between the CDFs implied by the empirical distribution of the human responses ($P$). This distance can be minimized using simple non-linear programming techniques. The weights $\mathbf{w}^{*}$ corresponding to the optimal mixture appear in the second column of Table 2.141414See https://www.expectedparrot.com/content/6f58d11f-98cc-4de5-bb89-edcf78042d79 for the agents.
@@ 3.3 Validation using game variants
To validate these prompts, we elicit their response distributions to the costless and cycle versions of the game. Both games maintain the fundamental requirement of similar but distinct data-generating processes. They involve reasoning well-explained by level-$k$ thinking, but feature payoff structures and incentives that differ from those of the basic game. For these human response distributions (previously shown in Table 1), the forward KL-divergence between the basic and costless game is $1.26$, and between the basic and cycle game is $1.00$. These divergences are substantial: each is more than three times larger than the divergence between the human and optimized agent distributions in the right panel of Figure 1. Accurate predictions on these distinct games would indicate that our optimized AI agents capture generalizable patterns rather than merely replicating the original training scenario.
@@
Notes: This figure displays empirical PMFs for the costless (top row) and cycle (bottom row) variants of the 11-20 game. The columns correspond to the Baseline (red), the Optimized AI agents (blue). Within each panel, the empirical PMFs from Arad and Rubinstein are imposed in black. The KL-divergence between each human and the AI response distribution is displayed in each panel. For both variants of the game, the selected AI agents (blue) are far closer to the human distribution than the baseline (red), even though the selected AI agents are constructed using only the basic version of the game.
Consistent with the basic version of the game and Gao et al. (2024), the baseline AI overwhelmingly selects 19 shekels in both variants, a considerable divergence from AR’s human subjects. In contrast, $\bm{\theta}^{*}$ is a far better predictor of both validation settings. The costless game in particular shows substantial improvement, with the KL-divergence decreasing by 84% ($d(P,\hat{P}_{\bm{\theta}^{*}})=0.15$ vs. baseline $d(P,\hat{P}_{0})=0.93$). It is almost perfectly predictive of the human responses when comparing the empirical PMFs. In the cycle game, the KL-divergence between the optimized AI and human responses is also reduced substantially, by 71% relative to the baseline ($d(P,\hat{P}_{\bm{\theta}^{*}})=0.28$ vs. $d(P,\hat{P}_{0})=0.95$). Taken together, these results are interpreted as evidence that $\bm{\theta}^{*}$ has effectively generalized to the validation data—data that was not used to construct its mixture. We therefore gain confidence that these agents may better predict entirely new settings that call for similar strategic reasoning.
@@ 3.4 Optimizing among atheoretical prompts
We now generate sets of arbitrary, atheoretical AI agents using the basic version of the game, which ultimately fail validation on the cycle and costless versions. These agents will offer a substantial contrast with $\bm{\theta}^{*}$ when applied to entirely novel games in the next subsection. This exercise also highlights two potential pitfalls of AI simulations addressed by our approach: (i) that without theoretically motivated candidate prompts, optimization may fail to even improve predictive power in-sample over the baseline, (ii) that atheoretical candidate sets can be optimized to effectively match particular samples of human data—even when such samples are obviously overfitting. The former is addressed by using samples of AI agents grounded in plausible theoretical mechanisms, and the latter is addressed by using training and testing data from distinct settings (or training and testing both across many different settings).
@@
Figure 3 shows that these two selected samples do not even offer a good in-sample fit. Each row corresponds to a different variant of the game—the top row is the basic. The columns represent different AI agent types, with the empirical PMFs from AR superimposed in black. The KL-divergence between the distributions in each panel is shown in the top left of each panel. After optimization, the in-sample KL-divergence between the selected AI agents and the humans in AR is $d(P,\hat{P}_{\bm{\theta}^{*}_{Hist}})=2.16$ and $d(P,\hat{P}_{\bm{\theta}^{*}_{MB}})=2.36$ for the historical figures and Myers-Briggs, respectively. These are not much better than the baseline $d(P,\hat{P}_{0})=2.7$ and far worse than the strategic AI agents $d(P,\hat{P}_{\bm{\theta}^{*}})=0.3$.
@@
Notes: (a) shows empirical PMFs for three variants of the 11-20 game from Arad and Rubinstein, compared to selected atheoretical AI agent samples optimized using only the basic version. Rows correspond to game variants, and columns correspond to AI agent types, with human data superimposed in black. Historical figures and Myers-Briggs subjects poorly match human distributions across all variants. The “Always Pick ‘N”’ set matches human data perfectly in-sample but fails to generalize. (b) shows KL-divergence between human and AI responses for games from Arad and Rubinstein. The lowest KL-divergence in each panel is indicated by a checkmark. Only the strategically-selected AI agents consistently improve over the baseline in all games.
@@
Figure 3 succinctly compares these results with the optimized mixture of strategic AI agents and the baseline, marking the best-performing sample in each setting. Only $\bm{\theta}^{*}$ consistently outperforms the baseline and generalizes across all settings. All three atheoretical samples are strictly worse than the baseline on both validation games.

### v4

@@ 2.3 Theory as guide towards generalizability
Economic theories—and theories from other social sciences that embrace methodological individualism (Friedman, 1953)—are well-suited to be cast into prompts, since good theory rests on the choices and actions of individuals. Specifically, we consider a prompt to be “theoretically grounded” when it instructs the LLM to make predictions about how a human would respond based on some underlying theory or model of human behavior. For example, a utility function that trades off one’s own payoff against inequality might become “You value your own earnings but dislike outcomes where you earn far more (or less) than others.” A highly capable and tool-using LLM could even be given that utility function and parameters directly (e.g., $u(x_{self},x_{other})=x_{self}-|x_{self}-x_{other}|$). A prospect-theoretic model reasonably maps to “You are risk averse in gains, risk seeking in losses, and probability weight the very likely and very unlikely outcomes”—and of course these could be parameterized as well.
This definition is necessarily broad because what constitutes a theory is also broad. One might consider good theories to be those that are parsimonious, predictive, and interpretable. We view these as effective guides for constructing prompts that are grounded in theory. But just as there is no universal rule as to what makes a theory good, there is not a singular all-encompassing playbook for candidate prompts. Here, the researcher must leverage their expertise appropriately to determine which theories are most relevant to the domain where predictions are desired. This should be relatively straightforward in practice. One can even simply ask an LLM to construct a candidate set of prompts based on a given paper and then adjust the prompts accordingly. For example, the linked notebook—which took 15 minutes to code—contains an example where we take PDFs of 5 well-known papers in the behavioral economics literature and generate 10 agents from each. [^9] https://expectedparrot.com/content/32751dae-7a69-430d-9f1a-0c3f50ccef5b
It is worth highlighting just how flexible such agents are as predictive models. They can make predictions in response to any setting described in natural language. This is usually impossible for traditional machine learning and mathematical economic models, which operate within a fixed parameter space. One cannot introduce additional covariates to a regression model once it is trained. Similarly, a prospect-theoretic model for evaluating risky gambles cannot instantaneously incorporate other contextually relevant factors—the recent performance of financial markets, the demographic composition of the experimental sample, or even something as mundane as whether participants are making decisions before or after lunch. AI agents, by contrast, can interpolate over such details through natural language instructions. Whether these agents actually leverage this flexibility effectively is, of course, an empirical question, but one that is easily testable.
@@ 2.4 A brief summary of the approach
    Optimize prompts on Training Data. Optimize the candidate prompts to best match the training data. This might involve selecting a mixture of prompts or adjusting trait parameters to minimize some statistical distance from observed responses. Confirm that the optimized sample outperforms the baseline LLM off-the-shelf on the training data.
@@ 2.5 Methods for optimizing prompts in-sample
We briefly describe two methods for optimizing prompts with a given set of training data. The first, the selection method, assumes we have a finite library of candidate prompts and selects (or mixes) them to best fit the training data. We apply this method to the experiments presented in Section 3 with data from Arad and Rubinstein (2012). An early version of this idea was suggested by Horton (2023), and several others have since explored applications (Leng et al., 2024; Xie et al., 2025; Bui et al., 2025). The second, the construction method, parameterizes a prompt template with numeric trait dimensions and optimizes those parameters. To the best of our knowledge, this method is novel. An application of the method is presented in Appendix A using data from Charness and Rabin (2002).
@@
A finite set of unique candidate natural language prompts is first specified. For each prompt $\theta\in\Theta$, the LLM is used to generate a predictive distribution $\hat{P}_{\theta}$. Let $P$ represent the observed ground-truth human distribution. The objective is to solve
@@
Alternatively, a prompt template can be parameterized by numeric trait variables. This is best illustrated with an example. Suppose $\phi_{1}$ and $\phi_{2}$ capture degrees of self-interest and inequity aversion, respectively. The prompt could be:
@@
$\hat{P}_{\theta}$ denotes the distribution induced by the LLM under parameter vector $\theta=(\phi_{1},\phi_{2})$. Given an observed human distribution $P$, the optimal parameters $\bm{\theta^{\star}}$ are found by solving $\min_{\theta}\;d\bigl(P,\hat{P}_{\theta}\bigr)$. In practice, this can be solved using any derivative-free optimization algorithm, such as Bayesian optimization or evolutionary algorithms. Note that one need not be limited to a single prompt; optimization can be solved using multiple templates or sets of agents with different values for the same template.
@@
To assess whether $\bm{\theta^{\star}}$ generalizes, we compare predictive fit against the baseline $\hat{P}_{0}$. This is an appealing reference measure because many distance metrics between distributions (like the KL divergence) are not easily interpretable. Furthermore, improvement over the baseline provides direct evidence that designing and optimizing prompts is a worthwhile endeavor in the first place. [^12] Strong baseline performance would simply reflect the LLM’s already high off-the-shelf predictive capacity.
Relative improvement can be computed as a difference-in-distances: $d\bigl(P,\hat{P}_{0}\bigr)\;-\;d\bigl(P,\hat{P}_{\bm{\theta^{\star}}}\bigr)$. A positive difference indicates that the prompts provide better predictive power than the baseline. When we optimize over a set of candidate prompts, we seek those that yield this positive difference on both the training and testing data—that would suggest that the set generalizes.
@@ 3 Predicting behavior in novel strategic games
We begin by briefly reviewing the level-$k$ model of strategic reasoning that originally motivated AR. This model provides plausible theoretical underpinnings linking the training, testing, and novel games in this section. We then describe the structure of the original 11-20 money request game and outline our procedure for constructing theory-grounded AI agents: optimizing their parameters on human data from AR’s original experiment, and validating their predictive performance on distinct but related variants also studied by AR. Finally, we introduce an entirely new set of games adapted from AR’s original design but featuring distinct numeric ranges and a novel participant sample recruited from Prolific. Agents are evaluated on their ability to predict human behavior in these never-before-seen games.
Although this section focuses on strategic reasoning games, the same procedure can be applied in any setting. In Appendix A, we replicate the process using the allocation games from Charness and Rabin (2002), which explore social preferences. Two other important differences in that section are that we: (i) optimize across multiple related settings rather than a single training setting, and (ii) utilize the construction method to build agents. We also present an additional preregistered experiment with human subjects to validate these agents in new settings.
@@ 3.1 Arad and Rubinstein (2012)’s 11-20 money request game
The level-$k$ model posits that players differ in how many steps ahead they consider when forming their strategies (Nagel, 1995; Stahl and Wilson, 1994, 1995). The model defines different types of players, from level-0 to level-$k$. Level-0 players use some predefined arbitrary decision rule, while level-$k$ players ($k\geq 1$) best respond assuming others are level-(k1)(k-1) reasoners. Such a model highlights the idea that players’ behavior depends not only on their decisions but also on their beliefs about how other people think.
@@
Interestingly, the 11-20 game does not have a pure strategy Nash equilibrium. The best response to any choice greater than 11 is to undercut the opponent by 1. But if the other player chooses 11, also selecting 11 is strictly dominated by every other strategy. The top row of Table 1 shows the unique symmetric mixed strategy Nash equilibrium for the game, with most of its density lying between 15 and 17 (levels 3 to 5).
@@
Finally, Table 1 provides results from two additional variants of the game from AR. Both variants of the game still involve two players selecting numbers between 11 and 20. They differ from the basic version in their payouts (see Appendix B for the full instructions). In the cycle version, players can earn a bonus of 20 shekels by undercutting the other’s request by exactly one or by selecting 20 when the other selects 11. This version has the same unique mixed strategy Nash Equilibrium as the basic game because the only affected strategy profiles are (20,11) and (11,20), which are outside the equilibrium support. However, if there is a distribution of level-$k$ thinkers—as opposed to perfectly rational game-theoretic best-responders—such a change might significantly impact play.
The costless version has the same bonus structure as the basic game, but with a different payout for choices below 20. Here, requesting 20 yields 20 shekels outright, while choosing a lower amount guarantees 17 shekels plus a 20-shekel bonus if the lower request is exactly one less than the other player’s. It is comparatively “costless” to continue undercutting. This game induces a symmetric mixed strategy Nash equilibrium that is more uniform across the choice set than the basic and cycle versions.
The human responses from these game variants, also from a similar sample of college students, are noticeably shifted towards 18-20 shekels—even with the basic and cycle sharing the same equilibrium. AR attributes this to the increased salience and payoff of selecting a higher number. AR concluded that the collective results from these three experiments are best explained by a mix of strategic types consisting of level-0, level-1, level-2, level-3, and random choosers.
@@ 3.2 Optimizing AI agents in-sample
We begin by eliciting the baseline AI’s response distribution ($\hat{P}_{0}$). We prompt Gpt-4o to play the basic version of the game 1,000 times without any additional instructions, setting the temperature to 1. Figure 1 displays the results. The left panel shows the empirical PMF of the baseline AI responses (red), along with the empirical human distribution $P$ from AR’s original experiment (vertical black lines). The baseline AI almost exclusively selects 19 shekels (87%), demonstrating limited variability. Using the forward KL divergence as $d(\cdot,\cdot)$ with the human distribution as the reference, the distance between these distributions is $d(P,\hat{P}_{0})=2.7$.
@@
Notes: This figure displays empirical PMFs for three samples playing the basic 11-20 money request game: human subjects from Arad and Rubinstein (the vertical black lines in both panels), the off-the-shelf baseline (left panel), and responses from our selected AI agents based on the weights (right panel). The KL divergence between the human and AI distributions is reported in the upper left corner of each panel.
@@
We elicit response distributions $\hat{P}_{\theta}$ for each candidate prompt $\theta\in\Theta_{Strategic}$ by prompting Gpt-4o 100 times per prompt. [^13] We then elicit each AI agent’s responses using Chain-of-Thought prompting, which encourages step-by-step reasoning before producing a final answer (Wei et al., 2024). We implement this through two sequential prompts. Prompt 1: {11-20 game instructions}. Reason out a few settings according to your personality and how others might respond. Prompt 2: {11-20 game instructions}. You previously had the following thoughts: {Response to prompt 1}. What amount of money would you request?. This procedure creates the agents (and their response distributions) over which we then optimize. We then employ the selection method to identify the optimal mixture of prompts that minimizes the absolute difference between the CDFs implied by the empirical distribution of the human responses ($P$). This distance can be minimized using simple nonlinear programming techniques. The weights $\mathbf{w}^{*}$ corresponding to the optimal mixture appear in the second column of Table 2.141414See https://www.expectedparrot.com/content/6f58d11f-98cc-4de5-bb89-edcf78042d79 for the agents.
@@ 3.3 Validation using game variants
To validate these agent sets, we elicit their response distributions to the costless and cycle versions of the game. Both games maintain our fundamental requirement of similar but distinct data-generating processes. They involve reasoning well-explained by level-$k$ thinking, but feature payoff structures and incentives that differ from those of the basic game. Most important, humans do not play these games in the same way as the basic version. The forward KL divergence between the basic and costless game is $1.26$, and between the basic and cycle game is $1.00$ (see Table 1 for the human response distributions for all three games). These divergences are substantial: each is more than three times larger than the divergence between the human and optimized agent distributions in the right panel of Figure 1. Accurate predictions on these distinct games would indicate that our optimized AI agents capture generalizable patterns rather than merely replicating the original training scenario.
@@
Notes: This figure displays empirical PMFs for the costless (top row) and cycle (bottom row) variants of the 11-20 game. The columns correspond to the baseline (red), the optimized AI agents (blue). Within each panel, the empirical PMFs from Arad and Rubinstein are imposed in black. The KL divergence between each human and the AI response distribution is displayed in each panel. For both variants of the game, the selected AI agents (blue) are far closer to the human distribution than the baseline (red), even though the selected AI agents are constructed using only the basic version of the game.
Consistent with the basic version of the game and Gao et al. (2024), the baseline AI overwhelmingly selects 19 shekels in both variants, a considerable divergence from AR’s human subjects. In contrast, $\bm{\theta}^{*}$ is a far better predictor of both validation settings. The costless game in particular shows substantial improvement, with the KL divergence decreasing by 84% ($d(P,\hat{P}_{\bm{\theta}^{*}})=0.15$ vs. baseline $d(P,\hat{P}_{0})=0.93$). It is almost perfectly predictive of the human responses when comparing the empirical PMFs. In the cycle game, the KL divergence between the optimized AI and human responses is also reduced substantially, by 71% relative to the baseline ($d(P,\hat{P}_{\bm{\theta}^{*}})=0.28$ vs. $d(P,\hat{P}_{0})=0.95$). Taken together, these results are interpreted as evidence that $\bm{\theta}^{*}$ has effectively generalized to the validation data—data that was not used to construct its mixture. We therefore gain confidence that these agents may better predict entirely new settings that call for similar strategic reasoning.
@@ 3.4 Optimizing among atheoretical prompts
We now generate sets of arbitrary, atheoretical AI agents using the basic version of the game, which ultimately fail validation on the cycle and costless versions. These agents will offer a substantial contrast with $\bm{\theta}^{*}$ when applied to entirely novel games in the next subsection. This exercise also highlights two potential pitfalls of AI simulations addressed by our approach: (i) without theoretically motivated candidate prompts, optimization may fail to even improve predictive power in-sample over the baseline, (ii) atheoretical candidate sets can be optimized to effectively match particular samples of human data—even when such samples are obviously overfitting. The former is addressed by using samples of AI agents grounded in plausible theory, and the latter is addressed by using training and testing data from distinct settings (or training and testing both across many different settings).
@@
Figure 3 shows that these two selected samples do not even offer a good in-sample fit. Each row corresponds to a different variant of the game—the top row is the basic. The columns represent different AI agent types, with the empirical PMFs from AR superimposed in black. The KL divergence between the distributions in each panel is shown in the top left of each panel. After optimization, the in-sample KL divergence between the selected AI agents and the humans in AR is $d(P,\hat{P}_{\bm{\theta}^{*}_{Hist}})=2.16$ and $d(P,\hat{P}_{\bm{\theta}^{*}_{MB}})=2.36$ for the historical figures and Myers-Briggs, respectively. These are not much better than the baseline $d(P,\hat{P}_{0})=2.7$ and far worse than the strategic AI agents $d(P,\hat{P}_{\bm{\theta}^{*}})=0.3$.
@@
Notes: (a) shows empirical PMFs for three variants of the 11-20 game from Arad and Rubinstein, compared to selected atheoretical AI agent samples optimized using only the basic version. Rows correspond to game variants, and columns correspond to AI agent types, with human data superimposed in black. Historical figures and Myers-Briggs subjects poorly match human distributions across all variants. The “Always Pick ‘N”’ set matches human data perfectly in-sample but fails to generalize. (b) shows KL divergence between human and AI responses for games from Arad and Rubinstein. The lowest KL divergence in each panel is indicated by a checkmark. Only the strategically-selected AI agents consistently improve over the baseline in all games.
@@
Figure 3 succinctly compares these results along with the optimized mixture of strategic AI agents and the baseline, marking the best-performing sample in each setting. Only $\bm{\theta}^{*}$ consistently outperforms the baseline and generalizes across all settings. All three atheoretical samples are strictly worse than the baseline on both validation games.